tax:

Algebra 2 check:

tax:

\[3(x - 4)^2 + 7 = 82\] Here was the equation

tax:

My final answer was: \[x = 9 , x = -1\]

tax:

@shadow Is this correct?

xXAikoXx:

@tax wrote:
My final answer was: \[x = 9 , x = -1\]
Valid, you isolated the squared term, divided by 3, took the square root, and solved both cases. \[3(x-3)^2 + 7 = 82\] \[3(x-4)^2=75\] \[(x-4)^2=25\] \[x-4=±5\] \[x - 4 = 5 ⇒ x - 9\] \[x−4=−5⇒x=−1\] 👍 you got it down

xXAikoXx:

@tax wrote:
@shadow Is this correct?
It's correct

tax:

@xxaikoxx wrote:
@tax wrote:
@shadow Is this correct?
It's correct
YAY Omgmg, thanks Aiko!